OpenStudy (jchick):

Daniel opened a car wash station in his neighborhood. He recorded the number of cars washed each week for the first six weeks. The data are shown as a scatter plot below. Scatter plot with line of best fit in the first quadrant titled Daniel's Car Wash Station, shows Week on the x axis and Number of Cars Washed on the y axis. The x axis scale is from 0 to 7 in increments of 1. The y axis scale is from 0 to 28 in increments of 4. The line of best fit passes through (0, 0), (3, 8).

OpenStudy (jchick):

Based on the line of best fit, which of these would be the best prediction for the number of cars washed in Week 9?

OpenStudy (littlebird):

options?

OpenStudy (jchick):

a. 8 b. 19 c. 24 d. 28

OpenStudy (littlebird):

I'm assuming since this has no picture, the problem wants you to use the coordinates provided to calculate an estimated equation for the line of best fit. In other words, use (0,0) and (3,8) to find y=mx+b

OpenStudy (jchick):

Here it is

OpenStudy (jchick):

My computer wouldn't upload it earlier

OpenStudy (littlebird):

The picture didn't change anything so..... m=slope which is (8-0)/(3-0) or 8/3 you find b by plugging in one of the coordinates 0=8/3(0) + b so... b=0 finally... y=8/3x where x represents the number of weeks. Plug in 9 for x and you get (8/3)x9=?

OpenStudy (jchick):

Idk how to do this

OpenStudy (jchick):

Oh wait I will do that sorry

OpenStudy (littlebird):

you solve these types of problems using formulas such as y=mx+b, m=(y1-y2)/(x1-x2), etc.

OpenStudy (jchick):

24 right?

OpenStudy (littlebird):

Yes

OpenStudy (jchick):

ok thanks!!!!